OpenStudy (anonymous):

What is the relationship between the lines determined by the following two equations? 15x – 3y = –12 y = 5x + 7 A. neither parallel nor perpendicular B. perpendicular C. parallel D. they are the same line

OpenStudy (anonymous):

@TheSmartOne I think I have a feeling of how to do this so hang on

TheSmartOne (thesmartone):

ok :)

OpenStudy (utterly_confuzzled):

Solve for y for the top equation -3y = -15x -12 y= 5x-12 They have the same slope, so that should give you the answer

OpenStudy (anonymous):

3y = -12 + 15x or 3y = 15x + -12 Then y = 15/3x - 4 @TheSmartOne is this correct?

OpenStudy (anonymous):

Then I would graph them.

TheSmartOne (thesmartone):

hold on

TheSmartOne (thesmartone):

it should be -3y for the first equation

TheSmartOne (thesmartone):

and if you're going to graph them at the end, you might as well graph them from what you were already given :P

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/eswtpvk3cm

OpenStudy (utterly_confuzzled):

But you don't have to graph them, just looking at the slopes should give you the answer

OpenStudy (anonymous):

Oh so would it be neither paralell or perpendicular?

OpenStudy (anonymous):

Or the same line?

TheSmartOne (thesmartone):

hold on @Abbster2015

TheSmartOne (thesmartone):

First let's make this 15x – 3y = –12 into slope-intercept form 15x – 3y = –12 We can rearrange the terms so: -3y + 15x = -12 subtract 15x on both sides :)

OpenStudy (anonymous):

-3y = 15x +12

TheSmartOne (thesmartone):

we need to subtract 15x, not add it :)

TheSmartOne (thesmartone):

so try again :)

OpenStudy (anonymous):

-3y = 15x - 12

OpenStudy (utterly_confuzzled):

If the slopes are the exact same, but they have different y-intercepts, then they are parallel If the slopes are negative opposite reciprocals, then they are perpendicular ex 7 and 7 would be parrllel, but 7 and (-1/7) would be perpendicular

TheSmartOne (thesmartone):

still wrong :P look at it this way -3y + 15x = -12 Subtract 15x on both sides -3y + 15x - 15x = -12 - 15x -3y = -12 - 15x And if you rearrange the terms -3y = -15x - 12

TheSmartOne (thesmartone):

now divide -3 on both sides :)

OpenStudy (anonymous):

y = -15/3x - 4

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

wrong and wrong :P

OpenStudy (anonymous):

Dang it

TheSmartOne (thesmartone):

we need to divide -3 on both sides so what is -15/ -3 = ? what is -12/ -3 = ? since both numbers have negatives, we can cross it out :) -15/-3 = 15/3 = ? -12/-3 = 12/3 = ?

OpenStudy (anonymous):

4

TheSmartOne (thesmartone):

and what is 15/3 = ?

OpenStudy (anonymous):

5

TheSmartOne (thesmartone):

so the final equation is y = 5x + 4

OpenStudy (anonymous):

Oh okay so it would be Paralell?

TheSmartOne (thesmartone):

and our second equation which was given to us was: y = 5x + 7 so they both have the same slope of 5... what does that mean? Are they parallel, perpendicular, both, or none? :P

OpenStudy (anonymous):

parallell?

TheSmartOne (thesmartone):

correct :)

OpenStudy (anonymous):

Yes! I am almost done now!